OpenStudy (abbycross167):

Can someone please help me to answer this math question? Please? I will fan and give a medal... Please help me to simplify this Rational Expression - 4x+2 10x-1 2x-3 ----- + ------ + ------ 5x 20x 4x

OpenStudy (abbycross167):

@AmTran_Bus could you please try to help me with this question?

OpenStudy (amtran_bus):

Type it into wolfram alpha (google that). Cant right now sorry.

OpenStudy (abbycross167):

oh okay

OpenStudy (abbycross167):

@phi can you help me with this question please?

OpenStudy (phi):

I would use a common denominator of 20x

OpenStudy (abbycross167):

yes sir

OpenStudy (phi):

that means multiply the first fraction by 4/4

OpenStudy (abbycross167):

so multiply by 4? 4x+2 ----- * 4? 5x

OpenStudy (phi):

not by 4, by \[ \frac{4}{4} \]

OpenStudy (abbycross167):

oh ok

OpenStudy (abbycross167):

what do I do with the 4x+2? do I go ahead and add the 2 to the 4?

OpenStudy (phi):

the top becomes (4x+2)*4 that is 4 x's plus a 2 x+x+x+x+1+1 if we expand it out (of course we usually write that as 4x+2)

OpenStudy (phi):

4* (4x+2) means you have 4 of the (4x+2) people use the "distributive law" to multiply 4 times each "thing" inside the parens but its good if you have an idea that helps to remember what to do. think of the parens as a package with stuff inside it, and you have 4 of these packages if you have that idea, it makes sense that you have 4 of each thing inside or if we write it out 4(4x+2) = 4*4x + 4*2 = 16x+8

OpenStudy (abbycross167):

ok so 4x+2 4 16x+8 ----- * --- = ------- 5x 4 20x ?

OpenStudy (abbycross167):

@phi ?

OpenStudy (phi):

yes. Let's skip the middle fraction for now (it already has 20x in the bottom) any idea what to multiply the bottom of the last fraction to make it 20x ?

OpenStudy (abbycross167):

6?

OpenStudy (phi):

try it: 6*4x= ?

OpenStudy (abbycross167):

5 sorry

OpenStudy (phi):

yes, 5 looks good if we multiply the bottom by 5, then we have to multiply the top by the same thing i.e. 5 can you multiply the top by 5 ?

OpenStudy (abbycross167):

10x-15?

OpenStudy (phi):

yes. now all 3 fractions have the same bottoms that means we can add the "tops" and put that over 20x to get the answer so we add up 16x + 8 10x - 1 10x - 15 can you do that ? you add the numbers "normally" for the x terms: think like this: 16 x's plus 10 x's plus 10 more x's how many x's is that all together?

OpenStudy (abbycross167):

36x?

OpenStudy (phi):

yes. and the numbers 8 -1 -15 ?

OpenStudy (abbycross167):

8? or -8?

OpenStudy (phi):

-8. if numbers have the same sign like this: -1 -15 then you can add them up and make the answer negative: -16 you now have 8 - 16 if you have two numbers with different signs subtract big - small: 16-8 = 8 then make it the sign of the "big" number (-16). we get -8

OpenStudy (phi):

so the answer so far is \[ \frac{36x-8}{20x} \] finally notice we can "factor out" 4 from the top. (this is the opposite of distributing) \[ \frac{4(8x-2)}{20x} \] now divide top and bottom by 4 to simplify

OpenStudy (phi):

* \[ \frac{4(9x-2)}{20x}\]

OpenStudy (phi):

you should get \[ \frac{9x-2}{5x}\]

OpenStudy (abbycross167):

thank you so much!! I really understand how to add them now!